OpenStudy (anonymous):

Phillip is using logarithms to solve the equation 63x = 18. Which of the following equations would be equivalent to his original expression? 6 log 3x = log 18 3 log 6 = x log 18 x log 3 = 6 log 18 3x log 6 = log 18

zepdrix (zepdrix):

I don't understand the original expression. Is it supposed to be 6(3^x) = 18 ?

zepdrix (zepdrix):

Or (63)^x = 18 maybe?

OpenStudy (anonymous):

the 3 x is an exponent

zepdrix (zepdrix):

6^(3x) = 18 We can take the log of each side: log[ 6^(3x) ] = log 18

zepdrix (zepdrix):

There is a rule of logs that allows us to deal with powers. log[ a^b ] = b*log a

zepdrix (zepdrix):

See how we can use that rule to move the 3x?

OpenStudy (anonymous):

yes but i dont get it

zepdrix (zepdrix):

So using the rule, log[ 6^(3x) ] = log 18 allows us to bring our 3x down in front of the log as a coefficient 3x log 6 = log 18